OpenStudy (anonymous):

Find the equation of the quadratic function with zeros 10 and 12 and a vertex at (11, -2).

OpenStudy (ybarrap):

The equation we seek is $$ f(x)=(x-a)(x-b)-c $$ Using the knowns: $$ f(10)=0=(10-a)(10-b)-c\\ f(12)=0=(12-a)(12-b)-c\\ f(11)=2=(11-a)(11-b)-c $$ We have 3 equations and 3 unknowns. There are 2 options. 1 solution or no solutions. I found no solutions. If we didn't have the vertex requirement, then the solution would be f(x)=(x-10)(x+12). This parabola has vertex <11,-1>. Moving the parabola down so that vertex is at <11,-2> moves the zeros and hence there is no solution with vertex at this location while keeping the zeros at x=10 and at x=12. But moving the parabola down so that the